identify the solution(s) of the system of equations if any -3x-4y=2 8y=-6x-4
Answer by LinnW(1048)

-3x-4y=2
8y=-6x-4
Reformat 8y=-6x-4
add 6x to each side
6x + 8y = -4
So the two equations are
-3x-4y=2
6x+8y=-4
Multiply the top equation by -2 giving
6x+8y = -4
The two equations are equivalent so the solutions for x and y are
all real numbers.